A group of students got the following scores in a test: 6, 9, 12, 15, 18, and 21. Consider samples of size 3 that can be drawn from this population


1
Expert's answer
2022-05-27T00:23:39-0400

There are a total of 20 combinations

The samples are

6,9,12




"Mean =\frac{6+9+12}{3} = 9"

6,9,15




"Mean =\frac{6+9+15}{3} = 10"

6,9,18




"Mean =\frac{6+9+18}{3} = 11"

6,9,21




"Mean =\frac{6+9+21}{3} = 12"

6,12,15




"Mean =\frac{6+12+15}{3} = 11"

6,12,18




"Mean =\frac{6+12+18}{3} = 12"

6,12,21




"Mean =\frac{6+12+21}{3} = 13"

6,15,18




"Mean =\frac{6+15+18}{3} = 13"

6,15,21




"Mean =\frac{6+15+21}{3} = 14"

6,18,21




"Mean =\frac{6+18+21}{3} = 15"

9,12,15




"Mean =\frac{9+12+15}{3} = 12"

9,12,18




"Mean =\frac{9+12+18}{3} = 13"

9,12,21



"Mean =\frac{9+12+21}{3} = 14"

9,15,18




"Mean =\frac{9+15+18}{3} = 14"

9,15,21




"Mean =\frac{9+15+21}{3} = 15"

9,18,21




"Mean =\frac{9+18+21}{3} = 16"

12,15,18




"Mean =\frac{12+15+18}{3} = 15"

12,15,21




"Mean =\frac{12+15+21}{3} = 16"

12,18,21




"Mean =\frac{12+18+21}{3} = 17"

15,18,21




"Mean =\frac{15+18+21}{3} = 18"